# Investigating cannibalistic millisecond pulsar binaries using MESA : new constraints from pulsar spin and mass evolution

### v1

@@ 3.4 Huntsman spiders
Huntsman spiders are considered to comprise of a red giant and a pulsar in a widening orbit (Strader et al., 2015, 2019). They evolve from a case B RLO phase in LMXBs. In terms of companion masses, they are similar to RBs, with the difference being that their orbits are an order of magnitude wider than RBs. Looking at Figure 1, we see that the binary evolution tracks reproduce the huntsmen regime irrespective of the value of $f_{\rm pulsar}$. Contrary to RBs and BWs, with a high $f_{\rm pulsar}$ the orbits do not widen with increasing $f_{\rm pulsar}$. Figure 5 (left panel, with $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=2.6$ d, shown by blue curves) shows that wider orbits correspond to $f_{\rm pulsar}=0.0$. Since these binaries are undergoing case B RLO, the companions to the NS are giant stars that have developed a deep convective envelope. Convective envelopes expand rapidly on mass loss and drive a high mass-transfer rate. When the radio pulsar wind turns on and irradiates the companion, it further increases the instantaneous mass loss from the companion. The companion detaches from its Roche lobe, pausing the RLO phase. The resulting effect is a decreased expansion of the orbits.
@@
We also find more conservative mass-accretion rates better reproducing the only current huntsman system with an observed spin period (2.66 ms for the source 1FGL J1417.7–440; Camilo et al., 2016). Table 1 describes the various properties of this huntsman source, as well as the other huntsman known so far, 2FGL J0846.0+2820. Fig 7 shows the final spin period distributions of binary tracks that end up in the huntsmen regime, corresponding to the three $\beta$ values investigated in our study (0.0, 0.3, and 0.7). Compared to BWs and RBs, more conservative mass transfer ($\beta<0.3$) is required for a NS with an initial mass of $1.3\,{\rm M}_{\odot}$ to spin-up to 2.66 ms. With a $\beta=0.3$, only NSs with initial masses $\gtrsim 1.5\,{\rm M}_{\odot}$are able to reproduce 1FGL J1417.7–440.
Assuming $\beta=0.0$, and looking at the final mass distributions of the NSs in Figure 8, we see the maximum final NS masses that we can get in huntsmen spiders. We also compare our distributions to the estimated masses of the pulsars in 1FGL J1417.7–440 ($1.62^{+0.43}_{-0.17}$; Swihart et al., 2018) and 2FGL J0846.0+2820 ($1.96^{+0.41}_{-0.41}\,{\rm M}_{\odot}$; Swihart et al., 2017). Even if the binaries had $\beta=0.0$, for a huntsman binary to have a NS with mass greater than 2.0 M, it should have $M^{i}_{\rm NS}>1.3\,{\rm M}_{\odot}$ (with 1.3 M, the maximum mass reached is 1.85 M). Therefore, with $\beta>0.0$, the initial mass should be even higher to get a high mass NS at the end of RLO.
@@
Figure 10: Orbital period versus companion mass for diverging binary tracks from our simulated grids that have similar NS spin as the observed huntsman source 1FGL J1417.7–440 (2.66 ms) within a 20 % margin (dashed gray curves). $M_{\rm c}$ is the mass of the companion to the NS and $P_{\rm orb}$ is the orbital period. We further show the binary that best matches the sources in its orbital evolution in blue curves, having $\beta=0.0$ ,$f_{\rm pulsar}=0.1$, $M^{i}_{\rm c}=2.0\,{\rm M}_{\odot}$, $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, and $P^{i}_{\rm orb}=2.6$ d.
The observed and estimated properties of the two currently known huntsmen spiders are described in Table 1. The spin of 1FGL J1417.7–440 (2.66 ms) can be used to constrain its evolutionary history. To reproduce the observed spin and estimated current pulsar mass with an initially canonical NS, we need a fairly conservative RLO phase ($\beta\lesssim 0.3$; see Figures 7 and 8). We also use an upper limit of 2.926 ms and a lower limit of 2.394 ms on the final spin to better identify the matching binaries, to account for 20% uncertainty in our calculations for the spin. Though pulsar irradiation is not necessary to reproduce the orbital properties of huntsman spiders, since it does have an effect on the orbit, we compare the different values of $f_{\rm pulsar}$. In Figure 10 we show the binary tracks that satisfy our selection criteria. We also compare these to the estimated orbital properties of 1FGL J1417.7–440 (mass of $0.28^{+0.07}_{-0.03}\rm\,M\odot$ and orbital period 5.37 d; Swihart et al., 2018; Strader et al., 2015), and see what tracks pass through the observed properties. We find that the binary track that best matches the source’s observed orbital properties (within the errors) corresponds to $\beta=0.0$ and $f_{\rm pulsar}=0.1$, and with the initial parameters, $M^{i}_{\rm c}=2.0\,{\rm M}_{\odot}$, $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, and $P^{i}_{\rm orb}=2.6$ d. The final spin of the pulsar is 2.73 ms, which is within 3% of the observed spin. This binary ends as a He-WD with mass 0.25 ${\rm M}_{\odot}$  and a $1.73\,{\rm M}_{\odot}$ NS in an orbit of 6.46 d.
@@
As for the source, 2FGL J0846.0+2820, since we do not know its spin, there is a wider range of binary tracks which could be used to explain it and almost all $f_{\rm pulsar}$ values can reproduce its orbital parameters (see Figure 1). Compared to 1FGL J1417.7–440 (and other spider systems), 2FGL J0846.0+2820 has a non-negligible eccentricity in its orbit (0.06; Swihart et al., 2017). This contradicts the general assumption that these systems are old enough for tidal forces to have efficiently circularized the orbits (Verbunt & Phinney, 1995). This observed eccentricity could be the result of the interaction with a circumbinary disc (Antoniadis, 2014) or inefficient tides (such as in radiative envelopes, which would be the case for MS stars; Hurley et al., 2002).
@@
Although we can reproduce the orbital parameters observed, we also need to study which diverging LMXB/IMXB systems can spin-up NSs to MSP spins. Figure 9 shows the allowed initial parameter space for the spin-up of NSs in diverging LMXBs/IMXBs. The presented figure shows grids of initial pulsar masses of 1.3 and $2.0\,{\rm M}_{\odot}$, for $\beta=0.3$ and $f_{\rm pulsar}=0.0$. We present the figures with $\beta=0.3$ instead of 0.0 as we leave room for uncertainties during the RLO phase that were not accounted for in our simulations. The colorbar highlights the region of the initial parameter space where the NS was able to spin-up to $\leq 30\rm\,ms$. Not all diverging binaries will be able to spin-up to MSP range. As we go from a less massive NS to a more massive NS, the peak of the spin distribution (systems with the maximum spin-up due to accretion) moves to higher companion masses and lower orbital periods, from centered around $2.0\,{\rm M}_{\odot}$ and 2.5–3.5 d to being centered around $2.5\,{\rm M}_{\odot}$ and 1.6–3.0 d.
@@
The boundary of the huntsmen parameter space depends on the initial $q$ and $P^{i}_{\rm orb}$ of the binary (see Figure 9). The range of acceptable $q$ is almost constant for both NS masses and corresponds to regions of high accreted mass. If the companion to the NS is not massive enough, the mass-transfer rate will be too low for efficient NS spin-up. If it is too massive, the RLO phase is very short as the star evolves quick. Additionally, if the RLO phase is super-Eddington, accretion would be further limited. The top boundary of the parameter space is limited due to the evolutionary state of the companion. Wider orbits than $\sim 3.5$ d results in the companion filling its Roche lobe after it has developed a deep convective envelope (late case B RLO), making the RLO phase short. In Figure 9 (right panel, with $M^{i}_{\rm NS}=2.0\,{\rm M}_{\odot}$), we also see an isolated binary with $M^{i}_{\rm c}=4.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=1.8$ d that has a NS with MSP-like spin. The final binary is a wide He-WD/NS binary (final WD mass $0.31\,{\rm M}_{\odot}$ and final period 33 d) with a final NS mass of $2.26\,{\rm M}_{\odot}$. The neighboring binaries do not drive high enough mass-transfer to spin-up their NS and we do not find a similar binary with a lower initial NS mass.
3.5 Transitional MSPs
@@
There is also a sub-category of compact MSPs that have been observed, binaries that switch from the rotation-powered state (where the spin-down of the pulsar dominates the evolution) to accretion-powered state (mass-transfer from the companion dominates) and vice-versa (see review by Papitto & de Martino, 2022). These binaries are called transitional MSPs (tMSPs), and so far there are 3 observed, PSR J102347.67+003841.2 (Archibald et al., 2009), IGR J18245–2452 (Papitto et al., 2013), and XSS J12270-4859 (Bassa et al., 2014). The companion masses for all of them are in the range 0.15 to 0.2 ${\rm M}_{\odot}$  and orbital periods 4 to 11 hr (Papitto & de Martino, 2022, and references therein), which put them in the RB regime of the observed parameter space. Figure 5 shows the three tMSPs along with some of the tracks from our simulated binaries. Two of these lie the near-$P_{\rm bif}$ tracks for $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ (gray lines in the figure), where the tracks differentiate for $f_{\rm pulsar}=0.0$ and 0.5. The point of differentiation of the two tracks is the turning on of the pulsar wind, and it explains the transitioning of the MSP binaries from accretion-powered to rotation powered state. Some other factors that should be considered to explain tMSPs, are accretion disc instabilities (Lasota, 2001) which are not yet fully understood.
@@ 4.1 Spider binary formation
We can compare our simulated spiders to those from previous works. While we find the formation of RBs and BWs to be linked, Chen et al. (2013) found RBs and BWs to come from two distinct populations of LMXBs, depending on the irradiation efficiency. For BWs, they found $f_{\rm pulsar}\lesssim 0.08$ to be the required irradiation efficiency, while for RBs, it was $f_{\rm pulsar}\gtrsim 0.08$. They explored mainly case A RLO systems, $M^{i}_{\rm c}$ around $1.0\,{\rm M}_{\odot}$, $M^{i}_{\rm NS}$ around $1.4\,{\rm M}_{\odot}$, and $P^{i}_{\rm orb}$ from 0.08 to 1.4 d, and found increased orbital expansion with increasing $f_{\rm pulsar}$. Qualitatively, we also see similar behavior in our results (see Section 3.3). Particularly for BWs, looking at similar systems in Figure 1, for $f_{\rm pular}=0.0$ to 0.05, we cover the lower-left region of the observed BW space. The degree to which the orbit expands for $f_{\rm pular}\gtrsim 0.1$ is lesser than Chen et al. (2013).
@@
For RBs, Chen et al. (2013) found them to be produced from case A binaries ($\sim 1$ d) with strong irradiation ($f_{\rm pulsar}\gtrsim 0.08$). In our work, the majority of RBs are produced at much wider initial orbits (near-$P_{\rm bif}$ at RLO onset, starting at around 2.0 to 2.4 d at ZAMS), along with high irradiation. With $f_{\rm pulsar}=0.5$, we also see rapid expansion of the case A binaries (solid black curve and $M^{i}_{\rm c}=1.5$ in the right panel of Figure 5). However, this expansion in orbits is not enough to bring them completely into the RB regime. Benvenuto et al. (2014) and Benvenuto et al. (2015) studied the effect of X-rays on the companion surface, as well as pulsar wind irradiation. They found that all BWs should have passed through a RB stage previously. We find that while not all BWs passed through the RB region, most binary tracks passing RB tracks ended up in the BW region of the observed parameter space, and the BWs that passed through the RB region, were H-surface deficient and in narrow orbits. The number of observed BWs to RBs observed is similar (approximately 63% and 37%, respectively), which implies their evolutionary lifetimes are comparable. For both BWs and RBs in our simulated tracks, by the time the spider phase begins the binaries are $\gtrsim 10^{9}$ yr old, as most of the prior evolution happened during the MS phase, which is long for low-mass stars.
@@
The need for a persistent magnetic braking was also suggested by Ginzburg & Quataert (2020) and Ginzburg & Quataert (2021), where they found that magnetic braking must remain active even when the companion has developed a fully-convective structure and the pulsar wind has turned on. While the authors focused their works on BWs, they suggested that the pulsar wind is too weak to remove enough mass from the pulsar companions. They looked at companions with initial masses of $1.0\,{\rm M}_{\odot}$. Their conclusion is similar to our results, as we also find the binary tracks not reproducing all of the observed BWs for initial companion masses of $1.0\,{\rm M}_{\odot}$ (see $f_{\rm pulsar}=0.5$ in right panel of Figure 5). With increasing companion masses, we cover more of the observed BW parameter space. The MSP source PSR J1953+1844 (in the globular cluster M71) that was recently discovered, seems to be the evolutionary link between tidarrens and RBs (Pan et al., 2023). It has an orbital period of 53.3 min and an estimated companion mass of 0.07 M. Since it is associated with a globular cluster (and would hence have a lower metallicity) and our work seeks to study Galactic spiders, we do not include it in our observed sources catalogue. Still, Pan et al. (2023) suggested that this source was formed from a similar process as UCXBs (with $P^{i}_{\rm orb}$ values near $P_{\rm bif}$), which is similar to our results, albeit at different metallicities.
@@ 4.2 Accretion efficiency and non-conservative evolution
As we saw in Section 2.3, in the literature non-conservative mass-accretion efficiencies ($\beta\gtrsim 0.5$) are favoured since they can reproduce the LMXB observations. We find that in order to reproduce the observed properties of spider binaries and the observed spins, more conservative efficiencies are better suited ($\beta\lesssim 0.3$). We also find that with a $\beta$ of 0.7 (see Figure 2), we can reproduce the observed spin and estimated pulsar mass of PSR J1738+0333, 5.85 ms and $1.47^{+0.07}_{-0.06}\,{\rm M}_{\odot}$, respectively. An assumption of non-conservative mass accretion throughout the entire evolution is proxy for any instabilities that might be affecting accretion (Illarionov & Sunyaev, 1975; Pringle, 1981; van Paradijs, 1996, for instance, propeller effects or thermal-viscous instabilities). We include the propeller phase in the NS spin evolution (see Section 2.3.2), by removing 10% of the transferred mass due to the propeller action. Since this phase is highly uncertain even in its effectiveness, we assume a low efficiency of loss of angular momentum during this phase. We also assumed an full efficiency for transfer of angular momentum from the accreted matter during RLO (see Section 2.3.1).
As for accretion disc instabilities, we include a model that describes the presence of a accretion disc above a critical mass-transfer rate, below which an X-ray irradiated disc would be unstable (see Section 2.3). Since we have already included inefficient mass accretion due to these two processes, the mass-accretion efficiency ($\epsilon=1-\beta$) which is set to a fixed value during the entire binary evolution accounts for any additional causes of mass or angular momentum loss. It also accounts for any uncertainties in our models of inefficient accretion, since we use simplistic approximations. However, the binary may evolve differently due to a high value of fixed $\beta$ throughout compared to a low $\beta$ with restricted accretion at some instances. To test this, we ran a MESA  model without the X-ray irradiated accretion disc assumptions and with $\beta=0.7$ and $f_{\rm pulsar}=0.0$, with $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$, $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, and $P^{i}_{\rm orb}=1.0$ d. We compared it to runs with same initial parameters, with $\beta=0.3$ and the X-ray irradiated disc assumption. We found that without the restriction on accretion, the NS spun up to 0.67 ms, while with the restriction it spun up to 2.9 ms. Assuming a fixed $\beta$ might result in overestimation of highly spinning pulsars, as most of these systems undergo case A RLO that is a long-lasting phase with moderately high mass-transfer rate ($10^{-10}\,{\rm M}_{\odot}\rm yr^{-1}$ to $10^{-8}\,{\rm M}_{\odot}\rm yr^{-1}$). Despite this difference, our results do not contradict the LMXB observations supporting highly non-conservative behavior.

### v2

@@ 3.4 Huntsman spiders
Huntsman spiders are considered to comprise of a red giant and a pulsar in a widening orbit (Strader et al., 2015, 2019). They evolve from a case B RLO phase in LMXBs. In terms of companion masses, they are similar to RBs, with the difference being that their orbits are an order of magnitude wider than RBs. Looking at Figure 1, we see that the binary evolution tracks reproduce the huntsmen regime irrespective of the value of $f_{\rm pulsar}$. Contrary to RBs and BWs, with a high $f_{\rm pulsar}$ the orbits do not widen with increasing $f_{\rm pulsar}$. Figure 6 (left panel, with $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=2.6$ d, shown by blue curves) shows that wider orbits correspond to $f_{\rm pulsar}=0.0$. Since these binaries are undergoing case B RLO, the companions to the NS are giant stars that have developed a deep convective envelope. Convective envelopes expand rapidly on mass loss and drive a high mass-transfer rate. When the radio pulsar wind turns on and irradiates the companion, it further increases the instantaneous mass loss from the companion. The companion detaches from its Roche lobe, pausing the RLO phase. We see the shortened RLO phase in Figure 5 (bottom row). The resulting effect is a decreased expansion of the orbits.
@@
We also find more conservative mass-accretion rates better reproducing the only current huntsman system with an observed spin period (2.66 ms for the source 1FGL J1417.7–440; Camilo et al., 2016). Table 1 describes the various properties of this huntsman source, as well as the other huntsman known so far, 2FGL J0846.0+2820. Figure 8 shows the final spin period distributions versus the final NS masses of binary tracks that end up in the huntsmen regime, corresponding to the three $\beta$ values investigated in our study (0.0, 0.3, and 0.7). Compared to BWs and RBs, more conservative mass transfer ($\beta<0.3$) is required for an NS with an initial mass of $1.3\,{\rm M}_{\odot}$ to spin up to 2.66 ms. With a $\beta=0.7$, the minimum spin period reached is 5.83 ms. With a $\beta=0.3$, only NSs with initial masses $\gtrsim 1.5\,{\rm M}_{\odot}$ are able to reproduce the pulsar spin of 1FGL J1417.7–440. We also compare our simulations to the estimated masses of the pulsars in 1FGL J1417.7–440 ($1.62^{+0.43}_{-0.17}$; Swihart et al., 2018) and 2FGL J0846.0+2820 ($1.96^{+0.41}_{-0.41}\,{\rm M}_{\odot}$; Swihart et al., 2017). If the binaries had fully conservative RLO ($\beta=0.0$), for a huntsman to have an NS with mass greater than 2.0 ${\rm M}_{\odot}$, it should have $M^{i}_{\rm NS}>1.3\,{\rm M}_{\odot}$ (with 1.3 ${\rm M}_{\odot}$, the maximum mass reached is 1.85 ${\rm M}_{\odot}$). Therefore, a huntsman with a confirmed high mass pulsar would have started RLO with a pulsar heavier than the canonical mass of 1.3 ${\rm M}_{\odot}$.
@@
Figure 10: Orbital period versus companion mass for diverging binary tracks from our simulated grids that have a similar NS spin as the observed huntsman source 1FGL J1417.7–440 (2.66 ms) within a 20 % margin (dashed gray curves). The term $M_{\rm c}$ is the mass of the companion to the NS, and $P_{\rm orb}$ is the orbital period. We further show the binary that best matches the sources in its orbital evolution in blue curves, where $\beta=0.0$ ,$f_{\rm pulsar}=0.1$, $M^{i}_{\rm c}=2.0\,{\rm M}_{\odot}$, $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, and $P^{i}_{\rm orb}=2.6$ d.
[…]
@@
As for the source 2FGL J0846.0+2820, since we do not know its spin, there is a wider range of binary tracks that could be used to explain it, and almost all $f_{\rm pulsar}$ values can reproduce its orbital parameters (see Figure 1). Compared to 1FGL J1417.7–440 (and other spider systems), 2FGL J0846.0+2820 has a non-negligible eccentricity in its orbit (0.06; Swihart et al., 2017). This contradicts the general assumption that these systems are old enough for tidal forces to have efficiently circularized the orbits (Verbunt & Phinney, 1995). This observed eccentricity could be the result of the interaction with a circumbinary disk (Antoniadis, 2014) or inefficient tides (such as in radiative envelopes, which would be the case for MS stars; Hurley et al., 2002).
@@
Although we can reproduce the orbital parameters observed, we also need to study which diverging LMXB/IMXB systems can spin up NSs to MSP spins. Figure 9 shows the allowed initial parameter space for the spin-up of NSs in diverging LMXBs/IMXBs. The presented figure shows grids of initial pulsar masses of 1.3 and $2.0\,{\rm M}_{\odot}$, for $\beta=0.3$ and $f_{\rm pulsar}=0.0$. We present the figures with $\beta=0.3$ instead of 0.0 as we leave room for uncertainties during the RLO phase that were not accounted for in our simulations. The color bar highlights the region of the initial parameter space where the NS was able to spin up to $\leq 30\rm\,ms$. Not all diverging binaries will be able to spin up to MSP range. As we go from a less massive NS to a more massive NS, the peak of the spin distribution (systems with the maximum spin-up due to accretion) moves to higher companion masses and lower orbital periods, from centered around $2.0\,{\rm M}_{\odot}$ and 2.5–3.5 d to being centered around $2.5\,{\rm M}_{\odot}$ and 1.6–3.0 d.
@@
The boundary of the huntsmen parameter space depends on the initial $q$ and $P^{i}_{\rm orb}$ of the binary (see Figure 9). The range of acceptable $q$ is almost constant for both NS masses and corresponds to regions of high accreted mass. If the companion to the NS is not massive enough, the mass-transfer rate will be too low for efficient NS spin-up. If it is too massive, the RLO phase is very short as the star evolves quick. If the RLO phase is super-Eddington, accretion would be further limited. The top boundary of the parameter space is limited due to the evolutionary state of the companion. Wider orbits than $\sim 3.5$ d results in the companion filling its Roche lobe after it has developed a deep convective envelope (late case B RLO), making the RLO phase short. In Figure 9 (right panel, with $M^{i}_{\rm NS}=2.0\,{\rm M}_{\odot}$), we also see an isolated binary with $M^{i}_{\rm c}=4.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=1.8$ d that has an NS with MSP-like spin. The final binary is a wide He-WD/NS binary (final WD mass $0.31\,{\rm M}_{\odot}$ and final period 33 d) with a final NS mass of $2.26\,{\rm M}_{\odot}$. The neighboring binaries do not drive high enough mass-transfer to spin up their NS, and we did not find a similar binary with a lower initial NS mass.
3.5 Transitional millisecond pulsars
@@
There is also a sub-category of compact MSPs that have been observed, binaries that switch from the rotation-powered state (where the spin-down of the pulsar dominates the evolution) to accretion-powered state (mass-transfer from the companion dominates) and vice-versa (see review by Papitto & de Martino, 2022). These binaries are called transitional MSPs (tMSPs), and so far there are 3 observed, PSR J1023+0038 (Archibald et al., 2009), IGR J18245–2452 (Papitto et al., 2013), and XSS J12270-4859 (Bassa et al., 2014). The companion masses for all of them are in the range 0.15 to 0.2 ${\rm M}_{\odot}$  and orbital periods 4 to 11 hr (Papitto & de Martino, 2022, and references therein), which put them in the RB regime of the observed parameter space. Figure 6 shows the three tMSPs along with some of the tracks from our simulated binaries. Two of these lie the near-$P_{\rm bif}$ tracks for $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ (gray lines in the figure), where the tracks differentiate for $f_{\rm pulsar}=0.0$ and 0.5. The point of differentiation of the two tracks is the turning on of the pulsar wind, and it explains the transitioning of the MSP binaries from accretion-powered to rotation powered state. Some other factors that should be considered to explain tMSPs are accretion disk instabilities (Lasota, 2001), which are not yet fully understood.
@@ 4.1 Spider binary formation
[…]
@@
For RBs, Chen et al. (2013) found them to be produced from case A binaries ($\sim 1$ d) with strong irradiation ($f_{\rm pulsar}\gtrsim 0.08$). In our work, the majority of RBs are produced at much wider initial orbits (near-$P_{\rm bif}$ at RLO onset, starting at around 2.0 to 2.4 d at ZAMS), along with high irradiation. With $f_{\rm pulsar}=0.5$, we also see rapid expansion of the case A binaries (solid black curve and $M^{i}_{\rm c}=1.5$ in the right panel of Figure 6). However, this expansion in orbits is not enough to bring them completely into the RB regime. Benvenuto et al. (2014) and Benvenuto et al. (2015) studied the effect of X-rays on the companion surface, as well as pulsar wind irradiation. They found that all BWs should have passed through a RB stage previously. We find that while not all BWs passed through the RB region, most binary tracks passing RB tracks ended up in the BW region of the observed parameter space, and the BWs that passed through the RB region, were H-surface deficient and in narrow orbits. The number of observed BWs to RBs observed is similar (approximately 63% and 37%, respectively), which implies their evolutionary lifetimes are comparable. For the BWs and RBs in our simulated tracks, by the time the spider phase begins the binaries are $\gtrsim 10^{9}$ yr old, as most of the prior evolution happened during the MS phase, which is long for low-mass stars.
@@
The need for a persistent magnetic braking was also suggested by Ginzburg & Quataert (2020) and Ginzburg & Quataert (2021), where they found that magnetic braking must remain active even when the companion has developed a fully convective structure and the pulsar wind has turned on. While the authors focused their works on BWs, they suggested that the pulsar wind is too weak to remove enough mass from the pulsar companions. They looked at companions with initial masses of $1.0\,{\rm M}_{\odot}$. Their conclusion is similar to our results, as we also find the binary tracks not reproducing all of the observed BWs for initial companion masses of $1.0\,{\rm M}_{\odot}$ (see $f_{\rm pulsar}=0.5$ in right panel of Figure 6). With increasing companion masses, we cover more of the observed BW parameter space. The MSP source PSR J1953+1844 (in the globular cluster M71) that was recently discovered, seems to be the evolutionary link between tidarrens and RBs (Pan et al., 2023). It has an orbital period of 53.3 min and an estimated companion mass of 0.07 ${\rm M}_{\odot}$. Since it is associated with a globular cluster (and would hence have a lower metallicity) and our work seeks to study Galactic spiders, we do not include it in our observed sources catalog. Still, Pan et al. (2023) suggested that this source was formed from a similar process as UCXBs (with $P^{i}_{\rm orb}$ values near $P_{\rm bif}$), which is similar to our results, albeit at different metallicities.
@@ 4.2 Accretion efficiency and non-conservative evolution
As we saw in Section 2.3, in the literature non-conservative mass-accretion efficiencies ($\beta\gtrsim 0.5$) are favored since they can reproduce the LMXB observations. We find that in order to reproduce the observed properties of spider binaries and the observed spins, more conservative efficiencies are better suited ($\beta\lesssim 0.3$). An assumption of non-conservative mass accretion throughout the entire evolution is proxy for any instabilities that might be affecting accretion (Illarionov & Sunyaev, 1975; Pringle, 1981; van Paradijs, 1996, for instance, propeller effects or thermal-viscous instabilities). We include the propeller phase in the NS spin evolution (see Section 2.3.2) by removing 10% of the transferred mass due to the propeller action. Since this phase is highly uncertain even in its effectiveness, we assume a low efficiency of loss of angular momentum during this phase. We also assumed an full efficiency for transfer of angular momentum from the accreted matter during RLO (see Section 2.3.1). Our results are similar to those by Kar et al. (2024), who found that lower $\beta$ values are required to reproduce accreting MSPs.
As for accretion disk instabilities, we include a model that describes the presence of a accretion disk above a critical mass-transfer rate, below which an X-ray irradiated disk would be unstable (see Section 2.3). Since we have already included inefficient mass accretion due to these two processes, the mass-accretion efficiency ($\epsilon=1-\beta$) that is set to a fixed value during the entire binary evolution accounts for any additional causes of mass or angular momentum loss. It also accounts for any uncertainties in our models of inefficient accretion, since we use simplistic approximations. A binary would evolve similarly due to a high value of fixed $\beta$ compared to a low $\beta$ with restricted accretion at some instances. To test this, we ran MESA  models without the X-ray irradiated accretion disk assumptions and with $\beta=0.7$ and 0.3, and $f_{\rm pulsar}=0.0$. The initial properties were $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$, $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, and $P^{i}_{\rm orb}=1.0$ d. We compared it to runs with same initial parameters and with the X-ray irradiated disk assumption, and found no significant differences in the evolution.